# Reflected generalized BDSDEs driven by non-homogeneous Lévy processes and obstacle problems for stochastic integro-PDEs with nonlinear Neumann boundary conditions

### v2

@@ 1 Introduction
Motivated by this latter connection, Pardoux and Zhang [38] introduced a new class of BSDEs that includes an additional finite-variation term, given by an integral with respect to a continuous increasing process interpreted as the boundary local time of a diffusion. These are called generalized BSDEs. This class has proved to be a powerful tool for providing probabilistic representations of solutions to systems of parabolic and elliptic semilinear PDEs with nonlinear Neumann boundary conditions. A solution on a time horizon $T\in(0,+\infty)$ with terminal value $\xi$ and drivers $f(\omega,t,y,z)$ and $h(\omega,t,y)$ is a pair of processes $(Y_{t},Z_{t})_{t\leq T}$ satisfying
@@
    $\mathcal{F}_{t}\triangleq\mathcal{F}_{t}^{L}\otimes\mathcal{F}_{t,T}^{B}\vee\mathcal{N},\quad t\in[0,T],$
    $(B_{t})_{t\leq T}$ is a one-dimensional Brownian motion, $(L_{t})_{t\leq T}$ is a non-homogeneous Lévy process independent of $(B_{t})_{t\leq T}$, and $\mathcal{F}_{s,t}^{B}:=\sigma\{B_{r}-B_{s}:s\leq r\leq t\}$, while $\mathcal{F}_{t}^{L}:=\mathcal{F}_{0,t}^{L}$, and $\mathcal{N}$ is the collection of $\mathbb{P}$-null sets.
@@
The paper is organized as follows. In Section 2, we specify the stochastic basis, introduce notations, and prove auxiliary results used throughout the paper. In Section 3, we state our assumptions on the data $(\xi,f,h,\kappa,g,S)$ of the RGBDSDE-NL (1.5), present a modified Itô formula for processes involving forward–backward integrals, and establish existence and uniqueness under our general stochastic conditions. Finally, in Section 4, we provide a probabilistic interpretation for solutions to a class of stochastic integro-partial differential equations with nonlinear Neumann boundary conditions.
@@ 2 Preliminaries
Let $T\in(0,+\infty)$ be a fixed horizon time. Throughout this paper, $(B_{t})_{t\leq T}$ will denote a standard one-dimensional Brownian motion and $(L_{t})_{t\leq T}$ a non-homogeneous Lévy process independent of $(B_{t})_{t\leq T}$ defined on the complete probability spaces $(\Omega_{1},\mathcal{F}_{1},\mathbb{P}_{1})$ and $(\Omega_{2},\mathcal{F}_{2},\mathbb{P}_{2})$, respectively.
We will denote by $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F}=\{\mathcal{F}_{t}\}_{0\leq t\leq T})$ the complete probability space, where $\Omega=\Omega_{1}\times\Omega_{2}$ $\mathcal{F}=\mathcal{F}_{1}\otimes\mathcal{F}_{2}$, $\mathbb{P}=\mathbb{P}_{1}\otimes\mathbb{P}_{2}$ and $\mathcal{F}_{t}\triangleq\mathcal{F}_{t}^{L}\otimes\mathcal{F}_{t,T}^{B}\vee\mathcal{N}$, where for any process $(\eta_{t})_{t\leq T}$; $\mathcal{F}_{s,t}^{\eta}=\sigma\{\eta_{r}-\eta_{s};\;s\leq r\leq t\}$, $\mathcal{F}_{t}^{\eta}:=\mathcal{F}_{0,t}^{\eta}$, $\mathbb{F}^{\eta}:=(\mathcal{F}_{t}^{\eta})_{t\leq T}$ and $\mathcal{N}$ is the collection of $\mathbb{P}$-null sets of $\mathcal{F}$. Note that the family of $\sigma$-algebras $\{\mathcal{F}_{t}\}_{t\leq T}$ is neither increasing nor decreasing; in particular, it is not a filtration. Any random variable $\zeta(\omega_{1})$, $\omega_{1}\in\Omega_{1}$ (resp. $\eta(\omega_{2})$, $\omega_{2}\in\Omega_{2}$) is considered as a random variable on $\Omega$ via the following identification: $\zeta(\omega)=\zeta(\omega_{1})$ (resp. $\eta(\omega)=\eta(\omega_{2})$), $\omega\triangleq(\omega_{1},\omega_{2})$.
By $\mathcal{B}(\mathbb{R})$, we mean the Borel $\sigma$-algebra of the real line $\mathbb{R}=(-\infty,+\infty)$. All stochastic processes are defined on the time interval $[0,T]$. Additionally, to simplify the notation, we omit the dependence on $\omega$ for any given process or random function. By convention, all brackets and stochastic integrals are assumed to be zero at time zero.
@@
An $\mathbb{F}^{L}$-adapted $\mathbb{R}$-valued process $L:=(L_{t})_{t\leq T}$ is a non-homogeneous Lévy process [^6] A non-homogeneous Lévy process $L$ can also be defined on a probability space without fixing a filtration a priori. In general, we work on a probability space equipped with the filtration generated by $L$; see [22, page 7] for details., if the following conditions hold:
@@
    $\mathbb{E}\left[e^{iuL_{t}}\right]=\exp\left\{\int_{0}^{t}\left(iub_{s}-\frac{c_{s}}{2}u^{2}+\int_{\mathbb{R}}\left(e^{iue}-1-iue\mathds{1}_{\{|e|\leq 1\}}\right)F_{s}(de)\right)ds\right\},$ (2.6)
    where $b:[0,T]\to\mathbb{R}$ and $c:[0,T]\to\mathbb{R}^{\ast,+}$ are Borel-measurable functions, and for each $s\in[0,T]$, $F_{s}$ is a measure on $\mathbb{R}$ that integrates $(1\wedge|e|^{2})$ and satisfies $F_{s}(\{0\})=0$. Moreover, for every $\mathsf{A}\in\mathcal{B}(\mathbb{R})$, the map $s\mapsto F_{s}(\mathsf{A})$ is Borel-measurable on $[0,T]$.
@@
In Definition 1, we did not impose regularity on the paths of $L$. However, since $L$ is an additive process (see Remark 2), it admits an RCLL modification (see [47, Ch 2. Theorem 11.5]). From now on, we consider this modification, ensuring that the left limits and jumps of $L$, denoted by $L_{t-}=\lim_{s\nearrow t}L_{s}$ and $\Delta L_{t}=L_{t}-L_{t-}$, are well-defined for all $t\in[0,T]$.
@@
    $\int_{0}^{T}\left(|b_{s}|+|c_{s}|+\int_{\mathbb{R}}\left(1\wedge|e|^{2}\right)F_{s}(de)\right)ds<+\infty$
@@
    There are positive constants $\epsilon,\mathfrak{C}>0$ such that
    $\int_{0}^{T}\int_{\{|e|>1\}}e^{ue}F_{s}(de)ds<+\infty,\quad\forall u\in\left[-(1+\epsilon)\mathfrak{C},(1+\epsilon)\mathfrak{C}\right].$
@@
    $\mathsf{b}_{t}=\int_{0}^{t}b_{s}\,ds,\quad\mathsf{c}_{t}=\int_{0}^{t}c_{s}\,ds,\quad\mathsf{F}_{t}(de)=\int_{0}^{t}F_{s}(de)\,ds.$
@@
  • Under assumptions (NL1) and (NL2), the process $L$ is a special $\mathbb{F}^{L}$-semimartingale with the canonical representation
    $L_{t}=\int_{0}^{t}\mathsf{b}_{s}\,ds+\int_{0}^{t}\sqrt{c_{s}}\,dW_{s}+\int_{0}^{t}\int_{\mathbb{R}}e\,\tilde{\mu}(ds,de).$ (2.7)
@@
    • $\mathsf{b}_{s}:=b_{s}+\int_{\mathbb{R}}e\mathds{1}_{\{\left|e\right|>1\}}F_{s}(de)$, with $b_{s}$ given in (2.6).
@@
      $\nu\left(\left[0,t\right]\times A\right):=\int_{0}^{t}\int_{A}F_{s}(de)\,ds,\quad A\in\mathcal{B}(\mathbb{R}).$ (2.8)
@@
Let $m_{t}^{(1)}=\mathbb{E}[L_{t}]$ and $m_{t}^{(i)}=\mathbb{E}[L^{i}_{t}]=\int_{0}^{t}\int_{-\infty}^{+\infty}e^{i}\nu(de,ds)$ for $i\geq 2$ and $t\in[0,T]$, with $\nu$ defined by (2.8). Note that $m^{(1)}_{t}$ is well-defined due to Lemma 3.
@@
With the non-homogeneous Lévy process $(L_{t})_{t\leq T}$, we associate the family of processes $(H^{(i)})_{i\geq 1}$ defined by $H_{t}^{(i)}=\sum_{k=1}^{i}\alpha_{i,k}X_{t}^{(k)}$ where the coefficients $\alpha_{i,j}$ correspond to the orthonormalization of the polynomials $1,e,e^{2},\cdots$ with respect to the measure
$\pi([0,t],de)=\int_{0}^{t}c_{s}\delta_{0}(de)ds+\int_{0}^{t}e^{2}F_{s}(de)ds.$
More precisely, the polynomials $q_{n}$ defined by $q_{n}(e)=\sum_{k=1}^{n}\alpha_{n,k}e^{k-1}$ are orthonormal with respect to the measure $\pi$, i.e.
$\int_{0}^{t}\int_{-\infty}^{+\infty}q_{n}(e)q_{m}(e)\pi(ds,de)=0~\mbox{ if }n\neq m.$
@@
$p_{n}(e)=eq_{n}(e)=\alpha_{n,n}e^{n}+\alpha_{n,n-1}e^{n-1}+\cdots+\alpha_{n,1}e.$
@@
$\langle H^{(i)},H^{(j)}\rangle_{t}$ $=$ $\sum_{k=1}^{i}\sum_{k^{\prime}=1}^{j}\alpha_{i,k}\alpha_{j,k^{\prime}}\left(\int_{0}^{t}\int_{-\infty}^{+\infty}e^{k+k^{\prime}}F_{s}(de)ds+\int_{0}^{t}c_{s}ds\mathds{1}_{\{k=k^{\prime}=1\}}\right)$
$=$ $\int_{0}^{t}\int_{-\infty}^{+\infty}\sum_{k=1}^{i}\sum_{k^{\prime}=1}^{j}\alpha_{i,k}\alpha_{j,k^{\prime}}e^{k+k^{\prime}-2}\left\{e^{2}F_{s}(de)\right\}ds+\alpha_{i,1}\alpha_{j,1}\int_{0}^{t}c_{s}ds$
$=$ $\int_{0}^{t}\int_{-\infty}^{+\infty}q_{i}(e)q_{j}(e)\pi(de,ds).$
@@
$\begin{split}[H^{(i)},X^{(j)}]_{t}&=\sum_{k=1}^{i}\alpha_{i,k}\left(\int_{0}^{t}c_{s}ds\mathds{1}_{\{k=j=1\}}+L^{(k+j)}_{t}\right)\\ &=\alpha_{i,i}L_{t}^{(i+j)}+\alpha_{i,i-1}L_{t}^{(i-1+j)}+\cdots+\alpha_{i,1}L_{t}^{(1+j)}+\alpha_{i,1}\int_{0}^{t}c_{s}ds\mathds{1}_{\{j=1\}}\end{split}$
@@ 3 RGBDSDE-NL with stochastic monotone coefficients
This section presents the first main result of the paper. We begin by stating the assumptions on the data $(\xi,f,h,\kappa,g,S)$ of the RGBDSDE-NL (1.5). We then provide an auxiliary result: an extension of the classical Itô formula, with a full description and proof. This extension yields key a priori estimates that are used throughout the paper to establish uniqueness and to streamline several calculations. Finally, under stochastic monotonicity and Lipschitz conditions on the drivers, we establish existence and uniqueness via Yosida regularization and Picard iteration.
@@ 3.1 Assumptions and technical results
    $\varphi_{t}$, $\phi_{t}$ and $\psi_{t}$ are $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$;
@@
In the following, let $\mathfrak{c}>0$ denotes a finite constant that may take different values from line to line. Additionally, we will use the notation $\mathfrak{c}_{\nu}$ to indicate that the constant $\mathfrak{c}$ depends on a specific set of parameters $\nu$. Finally, it should be noted that references to assumptions (H-M)-(i)-(ix)-(x) are equivalent to referencing (H-M)-(i), (H-M)-(ix) and (H-M)-(x) separately.
The following extension of the well-known Itô formula is crucial for the remainder of the paper, particularly in deriving useful a priori estimates between the solutions of the RGBDSDE-NL (1.5). The proof is analogous to the one provided in [36, Lemma 1.3], and is based on the techniques outlined in [42, Ch. II, Theorem 32].
@@
$\begin{split}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}&=\left|\mathcal{Y}_{0}\right|^{2}+\theta\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{0}\right|^{2}dV_{s}+\mu\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{0}\right|^{2}d\kappa_{s}+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{B}_{s}ds\\ &+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{C}_{s}d\kappa_{s}+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{G}_{s}\overleftarrow{dB}_{s}+2\sum_{k=1}^{d}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s-}\mathsf{Z}_{s}^{(k)}dH_{s}^{(k)}\\ &-\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathsf{G}_{s}\right|^{2}ds+\sum_{k,k^{\prime}=1}^{d}\int_{0}^{t}\mathsf{Z}_{s}^{(k)}\mathsf{Z}_{s}^{(k^{\prime})}\,d\big[H^{(k)},H^{(k^{\prime})}\big]_{s}+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}d\mathcal{K}_{s}.\end{split}$
@@
$\begin{split}\mathbb{E}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}&=\mathbb{E}\left|\mathcal{Y}_{0}\right|^{2}+\theta\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{s}\right|^{2}d\kappa_{s}+2\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{B}_{s}ds\\ &+2\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{C}_{s}d\kappa_{s}-\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathsf{G}_{s}\right|^{2}ds+\sum_{k=1}^{d}\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathsf{Z}_{s}^{(k)}\gamma^{(k)}_{s}\right|^{2}ds+\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}d\mathcal{K}_{s}.\end{split}$
@@
For the first part, it suffices to apply Lemma 6 with $\Psi(x,y,z)=e^{\theta x+\mu y}\left|z\right|^{2}$ and $\left(\mathsf{A}_{t}\right)_{t\leq T}=(V_{t},\kappa_{t})_{t\leq T}$. Following this, in the expectation calculation stage, we apply the Burkholder-Davis-Gundy inequality (see, e.g., [42, Ch IV. Theorem 48]), which implies

### v3

@@ 1 Introduction
Motivated by this latter connection, Pardoux and Zhang [37] introduced a new class of BSDEs that includes an additional finite-variation term, given by an integral with respect to a continuous increasing process interpreted as the boundary local time of a diffusion. These are called generalized BSDEs. This class has proved to be a powerful tool for providing probabilistic representations of solutions to systems of parabolic and elliptic semilinear PDEs. A solution on a time horizon $T\in(0,+\infty)$ with terminal value $\xi$ and drivers $f(\omega,t,y,z)$ and $h(\omega,t,y)$ is a pair of processes $(Y_{t},Z_{t})_{t\leq T}$ satisfying
@@
    $\mathcal{F}_{t}\triangleq\mathcal{F}_{t}^{L}\vee\mathcal{F}_{t,T}^{B},\quad t\in[0,T],$
    $(B_{t})_{t\leq T}$ is a one-dimensional Brownian motion, $(L_{t})_{t\leq T}$ is a non-homogeneous Lévy process independent of $(B_{t})_{t\leq T}$, and $\mathcal{F}_{s,t}^{B}=\sigma\{B_{r}-B_{s};\;s\leq r\leq t\}$ while $\mathcal{F}_{t}^{L}:=\mathcal{F}_{0,t}^{L}$ is augmented with null sets.
@@
The paper is organized as follows. In Section 2, we specify the stochastic basis, introduce notation, and prove auxiliary results used throughout the paper. In Section 3, we state our assumptions on the data $(\xi,f,h,\kappa,g,S)$ of the RGBDSDE-NL (1.5), present a modified Itô formula for processes involving forward–backward integrals, and establish existence and uniqueness under our general stochastic conditions. Finally, in Section 4, we provide a probabilistic interpretation for solutions to a class of stochastic integro-partial differential equations with nonlinear Neumann boundary conditions.
@@ 2 Preliminaries
Let $\mathbb{F}:=\left(\mathcal{F}_{t}\right)_{t\leq T}$ be a filtration defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ that satisfies the usual conditions of right-continuity and completeness. By $\mathcal{B}(\mathbb{R})$, we mean the Borel $\sigma$-algebra of the real line $\mathbb{R}=(-\infty,+\infty)$. All stochastic processes are defined on the time interval $[0,T]$. Additionally, to simplify the notation, we omit the dependence on $\omega$ for any given process or random function. By convention, all brackets and stochastic integrals are assumed to be zero at time zero.
@@
An $\mathbb{F}$-adapted $\mathbb{R}$-valued process $L:=(L_{t})_{t\leq T}$ is a non-homogeneous Lévy process, if the following conditions hold:
@@
    $\mathbb{E}\left[e^{iuL_{t}}\right]=\exp\left\{\int_{0}^{t}\left(iub_{s}-\frac{c_{s}}{2}u^{2}+\int_{\mathbb{R}}\left(e^{iuz}-1-iuz\mathds{1}_{\{|z|\leq 1\}}\right)F_{s}(dz)\right)ds\right\},$ (2.6)
    where $b:[0,T]\to\mathbb{R}$ and $c:[0,T]\to\mathbb{R}^{\ast,+}$ are Borel-measurable functions, and for each $s\in[0,T]$, $F_{s}$ is a measure on $\mathbb{R}$ that integrates $(1\wedge|z|^{2})$ and satisfies $F_{s}(\{0\})=0$. Moreover, for every $\mathsf{A}\in\mathcal{B}(\mathbb{R})$, the map $s\mapsto F_{s}(\mathsf{A})$ is Borel-measurable on $[0,T]$.
@@
In Definition 1, we did not impose regularity on the paths of $L$. However, since $L$ is an additive process (see Remark 2), it admits an RCLL modification (see [46, Ch 2. Theorem 11.5]). From now on, we consider this modification, ensuring that the left limits and jumps of $L$, denoted by $L_{t-}=\lim_{s\nearrow t}L_{s}$ and $\Delta L_{t}=L_{t}-L_{t-}$, are well-defined for all $t\in[0,T]$.
@@
    $\int_{0}^{T}\left(|b_{s}|+|c_{s}|+\int_{\mathbb{R}}\left(1\wedge|z|^{2}\right)F_{s}(dz)\right)ds<+\infty$
@@
    There are strictly positive constants $\epsilon$, $\mathfrak{C}$ such that
    $\int_{0}^{T}\int_{\{|z|>1\}}e^{uz}F_{s}(dz)ds<+\infty,\quad\forall u\in\left[-(1+\epsilon)\mathfrak{C},(1+\epsilon)\mathfrak{C}\right].$
@@
    $\mathsf{b}_{t}=\int_{0}^{t}b_{s}\,ds,\quad\mathsf{c}_{t}=\int_{0}^{t}c_{s}\,ds,\quad\mathsf{F}_{t}(dz)=\int_{0}^{t}F_{s}(dz)\,ds.$
@@
  • Under assumptions (NL1) and (NL2), the process $L$ is a special semimartingale with the canonical representation
    $L_{t}=\int_{0}^{t}\mathsf{b}_{s}\,ds+\int_{0}^{t}\sqrt{c_{s}}\,dW_{s}+\int_{0}^{t}\int_{\mathbb{R}}z\,\tilde{\mu}(ds,dz).$ (2.7)
@@
    • $\mathsf{b}_{s}:=b_{s}+\int_{\mathbb{R}}z\mathds{1}_{\{\left|z\right|>1\}}F_{s}(dz)$, with $b_{s}$ given in (2.6).
@@
      $\nu\left(\left[0,t\right]\times A\right):=\int_{0}^{t}\int_{A}F_{s}(dz)\,ds,\quad A\in\mathcal{B}(\mathbb{R}).$ (2.8)
Throughout this paper, $(B_{t})_{t\leq T}$ will denote a standard one-dimensional Brownian motion and $(L_{t})_{t\leq T}$ a non-homogeneous Lévy process independent of $(B_{t})_{t\leq T}$. Let $\mathcal{N}$ denotes the class of $\mathbb{P}$-null sets of $\mathcal{F}$. For each $t\in[0,T]$, we define
$\mathcal{F}_{t}\triangleq\mathcal{F}_{t}^{L}\vee\mathcal{F}_{t,T}^{B},\quad t\in[0,T],$
where for any process $(\eta_{t})_{t\leq T}$; $\mathcal{F}_{s,t}^{\eta}=\sigma\{\eta_{r}-\eta_{s};\;s\leq r\leq t\}\vee\mathcal{N}$, $\mathcal{F}_{t}^{\eta}:=\mathcal{F}_{0,t}^{\eta}$.
@@
Let $m_{t}^{(1)}=\mathbb{E}[L_{t}]$ and $m_{t}^{(i)}=\mathbb{E}[L^{i}_{t}]=\int_{0}^{t}\int_{-\infty}^{+\infty}x^{i}\nu(dx,ds)$ for $i\geq 2$ and $t\in[0,T]$, with $\nu$ defined by (2.8). Note that $m^{(1)}_{t}$ is well-defined due to Lemma 3.
@@
With the non-homogeneous Lévy process $(L_{t})_{t\leq T}$, we associate the family of processes $(H^{(i)})_{i\geq 1}$ defined by $H_{t}^{(i)}=\sum_{k=1}^{i}\alpha_{i,k}X_{t}^{(k)}$ where the coefficients $\alpha_{i,j}$ correspond to the orthonormalization of the polynomials $1,z,z^{2},\cdots$ with respect to the measure
$\pi([0,t],dz)=\int_{0}^{t}c_{s}\delta_{0}(dz)ds+\int_{0}^{t}z^{2}F_{s}(dz)ds.$
More precisely, the polynomials $q_{n}$ defined by $q_{n}(z)=\sum_{k=1}^{n}\alpha_{n,k}z^{k-1}$ are orthonormal with respect to the measure $\pi$, i.e.
$\int_{0}^{t}\int_{-\infty}^{+\infty}q_{n}(z)q_{m}(z)\pi(ds,dz)=0~\mbox{ if }n\neq m.$
@@
$p_{n}(z)=zq_{n}(z)=\alpha_{n,n}z^{n}+\alpha_{n,n-1}z^{n-1}+...+\alpha_{n,1}z.$
@@
$\langle H^{(i)},H^{(j)}\rangle_{t}$ $=$ $\sum_{k=1}^{i}\sum_{k^{\prime}=1}^{j}\alpha_{i,k}\alpha_{j,k^{\prime}}\left(\int_{0}^{t}\int_{-\infty}^{+\infty}z^{k+k^{\prime}}F_{s}(dz)ds+\int_{0}^{t}c_{s}ds\mathds{1}_{\{k=k^{\prime}=1\}}\right)$
$=$ $\int_{0}^{t}\int_{-\infty}^{+\infty}\sum_{k=1}^{i}\sum_{k^{\prime}=1}^{j}\alpha_{i,k}\alpha_{j,k^{\prime}}z^{k+k^{\prime}-2}\left\{z^{2}F_{s}(dz)\right\}ds+\alpha_{i,1}\alpha_{j,1}\int_{0}^{t}c_{s}ds$
$=$ $\int_{0}^{t}\int_{-\infty}^{+\infty}q_{i}(z)q_{j}(z)\pi(dz,ds).$
@@
$\begin{split}[H^{(i)},X^{(j)}]_{t}&=\sum_{k=1}^{i}\alpha_{i,k}\left(\int_{0}^{t}c_{s}ds\mathds{1}_{\{k=j=1\}}+L^{(k+j)}_{t}\right)\\ &=\alpha_{i,i}L_{t}^{(i+j)}+\alpha_{i,i-1}L_{t}^{(i-1+j)}+...+\alpha_{i,1}L_{t}^{(1+j)}+\alpha_{i,1}\int_{0}^{t}c_{s}ds\mathds{1}_{\{j=1\}}\end{split}$
@@ 3.1 Assumptions and technical results
    $\varphi_{t}$, $\phi_{t}$, and $\psi_{t}$ are $\mathcal{F}_{t}$-measurable for any $t\in[0,T]$;
@@
Remark 6.
The Skorokhod condition (1.5)-(iii) is equivalent to either $\int_{0}^{T}\left(Y_{s-}-S_{s}\right)dK_{s}=0$ or
$\int_{0}^{T}\left(Y_{s-}-S_{s-}\right)dK_{s}=0$. Indeed, since $S$ is continuous, we have $S_{s}=S_{s-}$. Moreover, as the reflection process $K$ is also continuous, the distinction between $Y_{s}$ and $Y_{s-}$ is irrelevant, because the set $\{s\in[0,T]:Y_{s-}\neq Y_{s}\}$ is finite (the number of jumps is d1d-1) and thus has $dK_{s}$-measure zero.
In the following, let $\mathfrak{c}>0$ denote a finite constant that may take different values from line to line. Additionally, we will use the notation $\mathfrak{c}_{\nu}$ to indicate that the constant $\mathfrak{c}$ depends on a specific set of parameters $\nu$. Finally, it should be noted that references to assumptions (H-M)-(i)-(ix)-(x) are equivalent to referencing (H-M)-(i), (H-M)-(ix) and (H-M)-(x) separately.
The following extension of the well-known Itô formula is crucial for the remainder of the paper, particularly in deriving useful a priori estimates between the solutions of the RGBDSDE-NL (1.5). The proof is analogous to the one provided in [35, Lemma 1.3], and is based on the techniques outlined in [41, Ch. II, Theorem 32].
@@
$\begin{split}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}&=\left|\mathcal{Y}_{0}\right|^{2}+\theta\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{0}\right|^{2}dV_{s}+\mu\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{0}\right|^{2}d\kappa_{s}+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{B}_{s}ds\\ &+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{C}_{s}d\kappa_{s}+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{G}_{s}\overleftarrow{dB}_{s}+2\sum_{k=1}^{d}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s-}\mathsf{Z}_{s}^{(k)}dH_{s}^{(k)}\\ &-\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathsf{G}_{s}\right|^{2}ds+\sum_{k,k^{\prime}=1}^{d}\int_{0}^{t}\mathsf{Z}_{s}^{(k)}\mathsf{Z}_{s}^{(k^{\prime})}\,d\big[H^{(k)},H^{(k^{\prime})}\big]_{s}+2\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}d\mathcal{K}_{s}\end{split}$
@@
$\begin{split}\mathbb{E}\Phi^{\theta,\mu}_{t}\left|\mathcal{Y}_{t}\right|^{2}&=\mathbb{E}\left|\mathcal{Y}_{0}\right|^{2}+\theta\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{s}\right|^{2}dV_{s}+\mu\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathcal{Y}_{s}\right|^{2}d\kappa_{s}+2\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{B}_{s}ds\\ &+2\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}\mathsf{C}_{s}d\kappa_{s}-\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathsf{G}_{s}\right|^{2}ds+\sum_{k=1}^{d}\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\left|\mathsf{Z}_{s}^{(k)}\gamma^{(k)}_{s}\right|^{2}ds+\mathbb{E}\int_{0}^{t}\Phi^{\theta,\mu}_{s}\mathcal{Y}_{s}d\mathcal{K}_{s}\end{split}$
@@
For the first part, it suffices to apply Lemma 7 with $\Psi(x,y,z)=e^{\theta x+\mu y}\left|z\right|^{2}$ and $\left(\mathsf{A}_{t}\right)_{t\leq T}=(V_{t},\kappa_{t})_{t\leq T}$. Following this, in the expectation calculation stage, we apply the Burkholder-Davis-Gundy (B-D-G) inequality (see, e.g., [41, Ch IV. Theorem 48]), which implies